2.1 Money demand in an economy in which no interest is paid on money is Y i P M d  1000  0.4 100 (a) Given that P = 100, Y = 1000, and i = 0.10. Find real money demand, nominal money demand, and velocity. (6) (b)The price level doubles from P =100 to P = 200. Find real money demand, nominal money demand, and velocity


1
Expert's answer
2021-05-19T10:56:46-0400

Solution:

Md = P * L (R, Y)

Where: P = Price level

            L = Liquidity function

            R = Interest rate

            Y = Real output


Md = 1000 + 0.4Y – 100i

a.). Real money demand:

Substitute the values in the demand function:

Md = 1000 + 0.4(1000) – 100(0.1)

Md = 1000 + 400 – 10

Md = 1390


Nominal money demand = Real money demand x Price

Nominal money demand = 1390 x 100 = 139,000


Money velocity (MV) = Price (P) x Real output (Y)

Money velocity (MV) = 100 x 1000 = 100,000


b.). An increase in P = 100 to P = 200:

Real money demand:

Substitute the values in the demand function:

Md = 1000 + 0.4(1000) – 100(0.1)

Md = 1000 + 400 – 10

Md = 1390


Nominal money demand = Real money demand x Price

Nominal money demand = 1390 x 200 = 278,000


Money velocity (MV) = Price (P) x Real output (Y)

Money velocity (MV) = 200 x 1000 = 200,000